# High-Performance Statistical Computing (HPSC): Challenges, Opportunities, and Future Directions

### v2

@@ 5.4 Energy-Efficient Statistical Computing
As HPC systems scale in processor count, memory capacity, and accelerator usage, energy consumption has become an increasingly pressing concern, both environmentally and economically. Large-scale statistical applications, including simulation-based inference, bootstrapping, and iterative optimization, often incur high computational costs that translate directly into significant energy usage. Designing energy-efficient statistical workflows is therefore critical, particularly as sustainability becomes a central consideration in scientific computing. One promising approach is mixed-precision computation, where select operations are executed in lower-precision formats (e.g., FP16 or bfloat16), reducing power consumption without compromising statistical accuracy. This strategy is especially effective on GPUs and AI accelerators optimized for low-precision arithmetic [Dongarra et al., 2024]. Mixed-precision methods have proven successful in accelerating iterative algorithms such as MLE, achieving faster convergence and reduced energy footprints [Cao et al., 2023].
Beyond precision tuning, algorithmic approximation techniques offer further gains in energy efficiency. Randomized numerical linear algebra methods—such as sketching, random projections, and low-rank matrix approximations—can significantly reduce computational load while maintaining acceptable accuracy [Baumann et al., 2024, Abdulah et al., 2018b]. Likewise, methods like approximate Bayesian computation (ABC) [Csilléry et al., 2010] and variational inference serve as scalable alternatives to traditional MCMC, enabling faster inference under limited resource budgets. These approximations are especially valuable in high-dimensional or streaming settings, where evaluating the whole model is often infeasible. As discussed throughout this review, such strategies are already making statistical computing more tractable across domains ranging from spatial modeling to genomics.
Hardware-aware optimization further amplifies these benefits. Adapting statistical algorithms to energy-efficient architectures—such as ARM processors [Ou et al., 2012] or vectorized instruction sets [Jakobs et al., 2016]—can yield substantial savings without sacrificing performance. Minimizing data movement is another key priority, as communication across memory hierarchies or networked systems often accounts for a significant portion of energy costs. Techniques such as communication-avoiding algorithms and asynchronous execution can mitigate these bottlenecks, enabling more sustainable computation at scale.
@@ 6.1 Statistical Computing on Specialized Hardware (GPU, TPU, IPU, DPUs)
Most existing R libraries and statistical software are optimized for CPU execution and offer limited support for GPU acceleration, typically through low-level C/C++ bindings. As datasets rapidly grow and models become more complex, this CPU-centric approach is increasingly insufficient. Tasks such as large-scale simulation, Bayesian inference, and high-dimensional modeling often exceed the capabilities of a single-node CPU, resulting in significant performance bottlenecks. Modern accelerator hardware, initially designed for machine learning, offers massive parallelism and high memory bandwidth that can dramatically speed up statistical workloads. GPUs, in particular, enable large-scale matrix operations and sampling routines to be executed orders of magnitude faster. By employing low-precision formats such as FP16 or BF16 where acceptable, and reserving FP32 or FP64 for numerically sensitive operations, modern platforms can cut memory usage, boost throughput, and reduce energy consumption [Ltaief et al., 2023, Cao et al., 2023]. However, statistical computing often requires high numerical accuracy, especially for tasks such as variance estimation, likelihood evaluation, and posterior sampling. Thus, future work must focus on adaptive mixed-precision strategies that dynamically adjust precision to preserve stability and avoid bias in intermediate computations [Abdulah et al., 2024a].
@@
The shift toward specialized hardware also opens the door to domain-specific accelerators. GPU tensor cores [Markidis et al., 2018], Intelligence Processing Units [IPUs, Louw and McIntosh-Smith, 2021], and Data Processing Units [DPUs, Barsellotti et al., 2022] offer compelling advantages for different classes of statistical computation. IPUs are well-suited for fine-grained parallelism in matrix-intensive tasks such as GPs and hierarchical Bayesian models. DPUs, on the other hand, can offload data movement and preprocessing operations—like shuffling, sorting, and filtering—freeing up the main processor for model training or inference. Looking further ahead, quantum computing presents a radically different paradigm. While still in its infancy, quantum accelerators have the potential to offer exponential speedups for certain statistical tasks [Yazdi, 2024]. Early exploration into quantum algorithms, such as quantum Monte Carlo methods, variational quantum eigensolvers [Tilly et al., 2022], and quantum-enhanced optimization—could unlock new approaches to Bayesian inference, MCMC, and high-dimensional clustering. Major challenges remain, including hardware instability, high error rates, and the need for hybrid quantum–classical workflows. However, investing in quantum circuit design, error correction techniques, and algorithmic robustness could now position the statistical computing community to fully leverage the benefits as quantum technologies mature. This is an achievable future direction for statisticians that requires understanding modern hardware architectures and optimizing existing statistical algorithms to fit these architectures well.
@@ 6.2 Federated Statistical Computing and Privacy-Preserving Inference
Initially developed in the machine learning community [McMahan et al., 2017], federated learning enables decentralized model training across multiple data sources without requiring raw data to be centrally aggregated. Building on this foundation, the emerging paradigm of Federated Statistical Computing (FSC) extends these principles to traditional statistical modeling. FSC provides a framework for performing inference across distributed datasets, including health records, financial transactions, and climate measurements, while respecting privacy constraints and regulatory boundaries [Jordan et al., 2019]. Instead of pooling raw data, federated statistical workflows compute local sufficient statistics, parameter estimates, or posterior approximations at each institution and securely aggregate them to form a global model. This preserves privacy while enabling reproducible, interpretable, and statistically valid inference. Active areas of research include the integration of privacy-preserving technologies such as differential privacy [Dwork and Lei, 2009] and homomorphic encryption [Lu et al., 2016], which provide strong formal guarantees for secure computation. Recent advances in federated Bayesian modeling [Zhang et al., 2022] and federated GPs [Yue and Kontar, 2024] demonstrate how complex probabilistic models can be trained without exposing sensitive data or intermediate computations.
@@
However, FSC introduces challenges distinct from those in federated learning. Statistical inference demands accurate uncertainty quantification, posterior consistency, and asymptotic guarantees—all of which must be revisited in settings characterized by communication bottlenecks, heterogeneous data distributions, and asynchronous updates across institutions. Novel algorithms, such as federated stochastic gradient Langevin dynamics [El Mekkaoui et al., 2021] and hierarchical model partitioning [Marion et al., 2018], aim to minimize communication overhead while maintaining statistical integrity. Applications of FSC are rapidly expanding into domains such as survival analysis, spatial modeling, and causal inference—especially in contexts where collaborative analysis is essential but centralized data collection is infeasible, as seen in healthcare, finance, and environmental science. As FSC evolves, it is poised to become a cornerstone of scalable, privacy-aware statistical computing. Future directions include developing federated analogs to classical statistical procedures, formalizing uncertainty-aware aggregation protocols, and integrating FSC into HPC workflows. In doing so, FSC closes a crucial gap between statistical methodology and the decentralized, privacy-constrained reality of modern data science. Figure 6.2 illustrates a typical FSC architecture, where multiple institutions compute and securely transmit local summaries—such as histograms, regression coefficients, or scatter plots—to a central server for privacy-preserving model integration. FSC is an achievable, low-risk direction. Recent studies are tackling privacy challenges, and the mature ML literature on federated learning offers clear guidelines for researchers.
[…]
@@
figureFederated Statistical Computing: Institutions share local statistical summaries with a central server to enable privacy-preserving model aggregation.
@@ 6.4 Novel Statistical Methods
One promising direction involves methods that are inherently parallel, rather than sequential algorithms awkwardly adapted to parallel settings. In scalable Bayesian inference, for instance, techniques like Consensus Monte Carlo [Scott, 2017] compute independent posterior samples across data subsets and combine them post hoc, enabling efficient inference across distributed datasets. Similarly, divide-and-conquer strategies for likelihood estimation and optimization [Zhang et al., 2013] partition computations into smaller, parallelizable subtasks, yielding dramatic runtime reductions for large-scale problems.
Approximation methods also play an important role. Methods such as variational inference, randomized sketching, and low-rank matrix approximations help to strike a balance between computational efficiency and statistical accuracy. Algorithms such as randomized singular value decomposition [Halko et al., 2011] and sketching-based regression [Woodruff et al., 2014] make it feasible to analyze massive datasets within memory and time constraints. Although approximations can introduce bias or affect uncertainty quantification, they are often essential to enable inference at scale. The growing convergence of statistical modeling and machine learning within HPC contexts is also driving hybrid methodologies that fuse the interpretability of statistical inference with the representational power of deep learning. Examples include deep Bayesian neural networks, probabilistic graphical models on GPUs, and statistical layers embedded within neural architectures—tools increasingly deployed in scientific domains where uncertainty and explainability are paramount.
At scale, communication costs become a significant bottleneck. Efficient statistical computing requires novel methods to manage this overhead, particularly for iterative or gradient-based algorithms. Communication-avoiding techniques, asynchronous updates, and decentralized learning strategies are actively being developed to mitigate these constraints [Ballard et al., 2014]. In tandem, hierarchical modeling frameworks that map naturally onto HPC hardware hierarchies—nodes, sockets, cores—can minimize costly inter-node data movement.
Emerging hardware platforms, such as GPUs, TPUs, and IPUs, offer new opportunities for designing statistical algorithms. Tools such as ExaGeoStat[Abdulah et al., 2018a] and pbdR[Ostrouchov et al., 2020] show how high-performance linear algebra and GPU acceleration can power scalable spatial and multivariate inference. Future research must continue to develop statistical kernels optimized for tensor cores, mixed-precision arithmetic, and hardware-specific instructions to ensure performance and accuracy on next-generation systems.
@@ 7 A Roadmap for the HPSC Community
Encouragingly, progress is already underway. Efforts such as the pbdR suite [Hasan et al., 2019], ExaGeoStat [Abdulah et al., 2018a], and other high-performance libraries demonstrate the viability of scalable statistical software that harnesses distributed-memory systems and accelerators. Interdisciplinary sessions at conferences such as Supercomputing (SC), Joint Statistical Meetings (JSM), and SIAM are also fostering dialogue between statisticians and computational scientists. However, these efforts remain fragmented—often limited to isolated projects without shared standards, cohesive tooling, or broad adoption. For HPSC to mature into a unified ecosystem, the community must move from isolated success stories to an integrated framework that supports sustainable growth and wide applicability.
@@
[…]
figureThe evolution of HPSC: transitioning from traditional, single-threaded R workflows to scalable, multi-node architectures empowered by parallel statistical algorithms, HPC programming models, and emerging technologies such as federated computing, quantum acceleration, and mixed-precision methods.

### v3

@@ 5.4 Energy-Efficient Statistical Computing
As HPC systems scale in processor count, memory capacity, and accelerator usage, energy consumption has become an increasingly pressing concern, both environmentally and economically. Large-scale statistical applications, including simulation-based inference, bootstrapping, and iterative optimization, often incur high computational costs that translate directly into significant energy usage. Designing energy-efficient statistical workflows is therefore critical, particularly as sustainability becomes a central consideration in scientific computing. One promising approach is mixed-precision computation, where select operations are executed in lower-precision formats (e.g., FP16 or bfloat16), reducing power consumption without compromising statistical accuracy. This strategy is especially effective on GPUs and AI accelerators optimized for low-precision arithmetic [73]. Mixed-precision methods have proven successful in accelerating iterative algorithms such as MLE, achieving faster convergence and reduced energy footprints [45].
Beyond precision tuning, algorithmic approximation techniques offer further gains in energy efficiency. Randomized numerical linear algebra methods—such as sketching, random projections, and low-rank matrix approximations—can significantly reduce computational load while maintaining acceptable accuracy [30, 10]. Likewise, methods like approximate Bayesian computation (ABC) [60] and variational inference serve as scalable alternatives to traditional MCMC, enabling faster inference under limited resource budgets. These approximations are especially valuable in high-dimensional or streaming settings, where evaluating the whole model is often infeasible. As discussed throughout this review, such strategies are already making statistical computing more tractable across domains ranging from spatial modeling to genomics.
Hardware-aware optimization further amplifies these benefits. Adapting statistical algorithms to energy-efficient architectures—such as ARM processors [198] or vectorized instruction sets [129]—can yield substantial savings without sacrificing performance. Minimizing data movement is another key priority, as communication across memory hierarchies or networked systems often accounts for a significant portion of energy costs. Techniques such as communication-avoiding algorithms and asynchronous execution can mitigate these bottlenecks, enabling more sustainable computation at scale.
@@ 6.1 Statistical Computing on Specialized Hardware (GPU, TPU, IPU, DPUs)
Most existing R libraries and statistical software are optimized for CPU execution and offer limited support for GPU acceleration, typically through low-level C/C++ bindings. As datasets rapidly grow and models become more complex, this CPU-centric approach is increasingly insufficient. Tasks such as large-scale simulation, Bayesian inference, and high-dimensional modeling often exceed the capabilities of a single-node CPU, resulting in significant performance bottlenecks. Modern accelerator hardware, initially designed for machine learning, offers massive parallelism and high memory bandwidth that can dramatically speed up statistical workloads. GPUs, in particular, enable large-scale matrix operations and sampling routines to be executed orders of magnitude faster. By employing low-precision formats such as FP16 or BF16 where acceptable, and reserving FP32 or FP64 for numerically sensitive operations, modern platforms can cut memory usage, boost throughput, and reduce energy consumption [165, 45]. However, statistical computing often requires high numerical accuracy, especially for tasks such as variance estimation, likelihood evaluation, and posterior sampling. Thus, future work must focus on adaptive mixed-precision strategies that dynamically adjust precision to preserve stability and avoid bias in intermediate computations [5].
@@
The shift toward specialized hardware also opens the door to domain-specific accelerators. GPU tensor cores [172], Intelligence Processing Units [IPUs, 162], and Data Processing Units [DPUs, 28] offer compelling advantages for different classes of statistical computation. IPUs are well-suited for fine-grained parallelism in matrix-intensive tasks such as GPs and hierarchical Bayesian models. DPUs, on the other hand, can offload data movement and preprocessing operations—like shuffling, sorting, and filtering—freeing up the main processor for model training or inference. Looking further ahead, quantum computing presents a radically different paradigm. While still in its infancy, quantum accelerators have the potential to offer exponential speedups for certain statistical tasks [275]. Early exploration into quantum algorithms, such as quantum Monte Carlo methods, variational quantum eigensolvers [251], and quantum-enhanced optimization—could unlock new approaches to Bayesian inference, MCMC, and high-dimensional clustering. Major challenges remain, including hardware instability, high error rates, and the need for hybrid quantum–classical workflows. However, investing in quantum circuit design, error correction techniques, and algorithmic robustness could now position the statistical computing community to fully leverage the benefits as quantum technologies mature. This is an achievable future direction for statisticians that requires understanding modern hardware architectures and optimizing existing statistical algorithms to fit these architectures well.
@@ 6.2 Federated Statistical Computing and Privacy-Preserving Inference
Initially developed in the machine learning community [173], federated learning enables decentralized model training across multiple data sources without requiring raw data to be centrally aggregated. Building on this foundation, the emerging paradigm of Federated Statistical Computing (FSC) extends these principles to traditional statistical modeling. FSC provides a framework for performing inference across distributed datasets, including health records, financial transactions, and climate measurements, while respecting privacy constraints and regulatory boundaries [134]. Instead of pooling raw data, federated statistical workflows compute local sufficient statistics, parameter estimates, or posterior approximations at each institution and securely aggregate them to form a global model. This preserves privacy while enabling reproducible, interpretable, and statistically valid inference. Active areas of research include the integration of privacy-preserving technologies such as differential privacy [76] and homomorphic encryption [166], which provide strong formal guarantees for secure computation. Recent advances in federated Bayesian modeling [284] and federated GPs [281] demonstrate how complex probabilistic models can be trained without exposing sensitive data or intermediate computations.
@@
However, FSC introduces challenges distinct from those in federated learning. Statistical inference demands accurate uncertainty quantification, posterior consistency, and asymptotic guarantees—all of which must be revisited in settings characterized by communication bottlenecks, heterogeneous data distributions, and asynchronous updates across institutions. Novel algorithms, such as federated stochastic gradient Langevin dynamics [77] and hierarchical model partitioning [171], aim to minimize communication overhead while maintaining statistical integrity. Applications of FSC are rapidly expanding into domains such as survival analysis, spatial modeling, and causal inference—especially in contexts where collaborative analysis is essential but centralized data collection is infeasible, as seen in healthcare, finance, and environmental science. As FSC evolves, it is poised to become a cornerstone of scalable, privacy-aware statistical computing. Future directions include developing federated analogs to classical statistical procedures, formalizing uncertainty-aware aggregation protocols, and integrating FSC into HPC workflows. In doing so, FSC closes a crucial gap between statistical methodology and the decentralized, privacy-constrained reality of modern data science. Figure 3 illustrates a typical FSC architecture, where multiple institutions compute and securely transmit local summaries—such as histograms, regression coefficients, or scatter plots—to a central server for privacy-preserving model integration. FSC is an achievable, low-risk direction. Recent studies are tackling privacy challenges, and the mature ML literature on federated learning offers clear guidelines for researchers.
@@
Figure 3: Federated Statistical Computing: Institutions share local statistical summaries with a central server to enable privacy-preserving model aggregation.
@@ 6.4 Novel Statistical Methods
One promising direction involves methods that are inherently parallel, rather than sequential algorithms awkwardly adapted to parallel settings. In scalable Bayesian inference, for instance, techniques like Consensus Monte Carlo [230] compute independent posterior samples across data subsets and combine them post hoc, enabling efficient inference across distributed datasets. Similarly, divide-and-conquer strategies for likelihood estimation and optimization [285] partition computations into smaller, parallelizable subtasks, yielding dramatic runtime reductions for large-scale problems.
Approximation methods also play an important role. Methods such as variational inference, randomized sketching, and low-rank matrix approximations help to strike a balance between computational efficiency and statistical accuracy. Algorithms such as randomized singular value decomposition [100] and sketching-based regression [265] make it feasible to analyze massive datasets within memory and time constraints. Although approximations can introduce bias or affect uncertainty quantification, they are often essential to enable inference at scale. The growing convergence of statistical modeling and machine learning within HPC contexts is also driving hybrid methodologies that fuse the interpretability of statistical inference with the representational power of deep learning. Examples include deep Bayesian neural networks, probabilistic graphical models on GPUs, and statistical layers embedded within neural architectures—tools increasingly deployed in scientific domains where uncertainty and explainability are paramount.
At scale, communication costs become a significant bottleneck. Efficient statistical computing requires novel methods to manage this overhead, particularly for iterative or gradient-based algorithms. Communication-avoiding techniques, asynchronous updates, and decentralized learning strategies are actively being developed to mitigate these constraints [25]. In tandem, hierarchical modeling frameworks that map naturally onto HPC hardware hierarchies—nodes, sockets, cores—can minimize costly inter-node data movement.
Emerging hardware platforms, such as GPUs, TPUs, and IPUs, offer new opportunities for designing statistical algorithms. Tools such as ExaGeoStat[9] and pbdR[197] show how high-performance linear algebra and GPU acceleration can power scalable spatial and multivariate inference. Future research must continue to develop statistical kernels optimized for tensor cores, mixed-precision arithmetic, and hardware-specific instructions to ensure performance and accuracy on next-generation systems.
@@ 7 A Roadmap for the HPSC Community
Encouragingly, progress is already underway. Efforts such as the pbdR suite [102], ExaGeoStat [9], and other high-performance libraries demonstrate the viability of scalable statistical software that harnesses distributed-memory systems and accelerators. Interdisciplinary sessions at conferences such as Supercomputing (SC), Joint Statistical Meetings (JSM), and SIAM are also fostering dialogue between statisticians and computational scientists. However, these efforts remain fragmented—often limited to isolated projects without shared standards, cohesive tooling, or broad adoption. For HPSC to mature into a unified ecosystem, the community must move from isolated success stories to an integrated framework that supports sustainable growth and wide applicability.
@@
Figure 4: The evolution of HPSC: transitioning from traditional, single-threaded R workflows to scalable, multi-node architectures empowered by parallel statistical algorithms, HPC programming models, and emerging technologies such as federated computing, quantum acceleration, and mixed-precision methods.